If I want to make the perfect 12 oz. cup of chocolate milk, it requires that the mixture is 42% Syrup. What I have now is 8 oz. of a milk/syrup mixture that I know contains 30% syrup. What must the syrup concentration be in the remaining mixture that I must add in order to achieve perfection?
Answer by stanbon(75887)

If I want to make the perfect 12 oz. cup of chocolate milk, it requires that the mixture is 42% Syrup. What I have now is 8 oz. of a milk/syrup mixture that I know contains 30% syrup. What must the syrup concentration be in the remaining mixture that I must add in order to achieve perfection?
-------------
Final mixture DATA:
Amt of mix = 12 oz. ; amt of syrup = 0.42(12) = 5.04 oz
---------------------------
Have now DATA:
Amt of mix = 8 oz. ; amt of syrup = 0.30(8) = 2.4 oz
------------------------------
Misture to be added DATA:
Amt of mix = 12-8 = 4 oz ; amt of syrup = 4x where x is the percent.
-------------------
EQUATION:
syrup + syrup = syrup
2.4 + 4x = 5.04
4x = 2.64x = .66
---------------
66% of the added mixture must be syrup
=============
Cheers,
Stan H.